# ComFairGNN: Community Fair Graph Neural Network

### v1

@@ 3.3. Training Objective and Constraints:
(11) $\mathcal{L}_{task}=-\frac{1}{|Y|}\sum_{i}\sum_{j}Y_{i,j}\log(\hat{Y}_{i,j})$
@@
(12) $\mathcal{L}_{fair}=\left|\frac{1}{|c_{y=0}|}\sum\limits_{i,j\in c_{y=0}}\text{sim}(x_{i},x_{j})-\frac{1}{|c_{y=1}|}\sum\limits_{i,j\in c_{y=1}}\text{sim}(x_{i},x_{j})\right|$
@@
(13) $\mathcal{L}=\mathcal{L}_{\text{task}}+\mathcal{L}_{\text{fair}}$
4. Experiments
@@
Dataset Comm-Id Comm(size) Group_0 (size) Group_1 (size)
Pokec-z 1 309 293 16
2 282 26 256
3 259 28 231
4 1412 1019 393
5 302 256 46
Pokec-n 1 428 407 21
2 360 58 302
3 361 43 318
3 641 28 231
4 389 322 67
Facebook 1 55 27 28
2 37 29 8
3 65 49 16
Table 1. Communities Statistics. Group_0 and Group_1 indicate the demographic subgroups.
@@
Table 2. Comparison of GNNs on utility and bias at the graph level with different debiasing methods.
[…]
Vanilla GCN 1 62.36 57.89 45.38 21.36 -45.38 -21.36
2 76.84 71.67 7.32 10.34 7.32 10.34
3 64.73 63.85 14.97 12.59 14.97 12.59
4 72.00 72.24 11.38 4.14 11.38 4.14
5 84.43 74.76 10.66 0.64 10.66 0.64
FairGNN 1 61.34 56.56 2.81 9.67 2.81 -9.67
2 75.71 71.58 7.23 7.97 -7.23 -7.97
3 61.43 60.26 21.19 18.21 21.19 18.21
4 69.14 69.36 11.76 13.29 -11.76 -13.29
5 79.24 66.35 13.18 6.22 13.18 6.22
NIFTY 1 61.56 56.46 1.31 4.69 1.31 -4.69
2 77.97 70.74 13.45 14.68 -13.45 -14.68
3 61.63 60.00 9.77 14.14 9.77 14.14
4 68.57 66.67 5.16 4.62 -5.16 -4.62
5 79.93 63.56 13.74 11.16 13.74 11.16
UGE 1 60.88 51.71 21.29 27.68 -21.29 -27.68
2 69.49 61.91 10.42 2.72 -10.42 -2.72
3 61.24 58.91 1.23 11.91 1.23 11.91
4 58.48 61.34 51.89 60.50 -51.89 -60.50
5 78.55 62.69 3.77 6.00 -3.77 -6.00
ComFairGNN 1 74.43 72.97 5.63 1.04 -5.63 -1.04
2 73.44 75.98 7.21 0.07 7.21 0.07
3 72.97 73.26 3.46 1.90 3.46 1.90
4 72.97 65.44 2.59 3.40 -2.59 -3.40
5 76.67 74.05 2.61 1.08 2.61 1.08
Table 3. Comparison of GNNs with debiasing methods on utility and bias for Pokec-z dataset at community level. The smaller $\Delta\textbf{EO}$ and $\Delta\textbf{SP}$ the better. $(w)$ and $(w/o)$ indicate with absolute value and without absolute value, respectively.
[…]
Vanilla GCN 1 72.30 73.73 15.20 6.72 -15.20 -6.72
2 69.19 68.00 13.06 2.74 13.06 2.74
3 66.11 59.05 26.67 14.32 -26.67 -14.32
4 71.39 71.96 8.29 6.16 8.29 6.16
5 68.42 66.25 19.14 8.86 19.14 8.86
FairGNN 1 65.68 64.42 22.65 12.27 22.65 12.27
2 63.17 62.71 9.67 9.43 9.67 9.43
3 58.68 55.46 16.35 17.85 -16.35 -17.85
4 74.89 62.42 10.60 8.57 10.60 8.57
5 60.23 57.35 4.20 4.76 -4.20 -4.76
NIFTY 1 72.97 72.19 23.17 29.61 23.17 29.61
2 67.23 66.10 4.29 1.54 4.29 1.54
3 66.39 58.59 9.44 1.59 -9.44 1.59
4 79.65 68.73 8.29 8.09 8.29 8.09
5 69.59 67.68 7.34 14.19 -7.34 -14.19
UGE 1 62.70 61.14 33.67 6.36 33.67 6.36
2 63.73 62.45 15.82 18.03 15.82 18.03
3 64.71 55.57 37.78 30.09 -37.78 -30.09
4 73.16 59.70 7.08 4.41 -7.08 -4.41
5 57.31 55.75 20.89 2.86 20.89 2.86
ComFairGNN 1 66.82 61.88 11.65 4.00 11.65 4.00
2 70.56 70.87 3.16 1.75 -3.16 -1.75
3 65.62 65.67 0.17 2.40 -0.17 -2.40
4 72.49 64.41 8.59 5.45 -8.59 -5.45
5 69.4 70.16 1.75 1.24 1.75 1.24
Table 4. Comparison of GNNs with debiasing methods on utility and bias for Pokec-n dataset at community level. The smaller $\Delta\textbf{EO}$ and $\Delta\textbf{SP}$ the better. $(w)$ and $(w/o)$ indicate with absolute value and without absolute value, respectively.
[…]
Vanilla GCN 1 83.02 54.42 5.26 8.33 5.26 8.33
2 72.97 68.59 23.53 21.98 -23.53 -21.98
3 85.07 54.55 1.49 4.55 .49 4.55
FairGNN 1 94.83 49.11 5.26 5.00 5.26 5.00
2 70.00 62.50 1.10 20.57 -1.10 20.57
3 84.06 57.46 5.00 6.96 5.00 6.96
NIFTY 1 81.82 50.00 0.00 0.00 0.00 0.00
2 67.57 56.68 24.68 29.31 24.68 29.31
3 78.46 52.59 2.56 4.21 -2.56 4.21
UGE 1 76.36 62.22 3.36 13.87 3.36 13.87
2 36.36 50.00 3.78 4.14 3.78 4.14
3 68.12 50.58 0.99 4.83 -0.99 4.83
ComFairGNN 1 81.78 77.12 2.43 4.75 2.43 4.75
2 63.41 52.61 2.51 1.76 -2.51 -1.76
3 81.13 50.00 0.09 0.01 0.09 0.01
Table 5. Comparison of GNNs with debiasing methods on utility and bias for Facebook dataset at community level. The smaller $\Delta\textbf{EO}$ and $\Delta\textbf{SP}$ the better. $(w)$ and $(w/o)$ indicate with absolute value and without absolute value, respectively.
4.1. Experimental Setup and Datasets
Datasets. In this context, the downstream task is node classification. We conducted our experiments using three real-world datasets: Pokec-z, Pokec-n, and Facebook. Specifically, Pokec-z and Pokec-n datasets are derived from Pokec, a well-known social network in Slovakia (Takac and Zabovsky, 2012). This dataset encompasses anonymized information from the entire network. Due to its extensive size, the authors in (Dai and Wang, 2021) extracted two subsets from the original Pokec dataset, designated as Pokec-z and Pokec-n, based on users’ provincial affiliations. In the dataset, each user is represented as a node, and each edge represents a friendship between two users. The sensitive attribute in this context is the users’ location. The objective is to classify users based on their field of work. The Facebook dataset is derived from the Facebook DataSet (He and McAuley, 2016). In this dataset, each user is represented as a node, and each edge signifies a friendship between two users. The sensitive attribute is the gender information associated with each node. We summarize the datasets in Table 6
Dataset Nodes Edges Features Classes
Pokec-z 67,796 882,765 276 2
Pokec-n 66,569 198,353 265 2
Facebook 1,034 26,749 573 2
Table 6. Summary of datasets.
Baseline Methods: We investigate the effectiveness of four state-of-the-art GNN debiasing methods at a community level, namely FairGNN (Dai and Wang, 2021), NIFTY (uNIfying Fairness and stabiliTY) (Agarwal et al., 2021a), and UGE (Wang et al., 2022a). FairGNN utilizes adversarial training to eliminate sensitive attribute information from node embeddings. NIFTY improves fairness by aligning predictions based on both perturbed and unperturbed sensitive attributes. This approach generates graph counterfactuals by inverting the sensitive feature values for all nodes while maintaining all other attributes. The method then incorporates a regularization term in the loss function to promote similarity between node embeddings derived from the original graph and its counterfactual counterpart. UGE derives node embeddings from a neutral graph that is free of biases and unaffected by sensitive node attributes, addressing the issue of unbiased graph embedding. We use GCN (Kipf and Welling, 2016) as the backbone GNN model.
Evaluation: We use the widely adopted group fairness metrics, $\Delta$SP (Statistical Parity), and $\Delta$EO (Equal Opportunity)   to evaluate fairness at a community level in Tables 345 and graph level in Table 2. As discussed in section 2.3, $\Delta$SP and $\Delta$EO fundamentally require that the model’s outcomes remain statistically independent of the nodes’ sensitive features. In particular, we form two groups of nodes from the test set, $S_{0}$ and $S_{1}$, containing test nodes with sensitive features 00 and $1$, respectively. The two groups fairness is the most serious fairness issue in GNN’s tasks. As GNNs learn by neighborhood aggregation section 2.1, we evaluate fairness within different subgraphs or communities within the larger graph. This can help identify and address local disparities more effectively. We select $S_{0}$ and $S_{1}$ nodes within the community to evaluate the model’s disparity locally, which may present more prominent biases. We select top and bottom $15$ sample homophily neighborhood distribution ratio from each community to formulate the representative coreset nodes from the training set during training. These nodes are sufficient to cover the local contextual structures where biases are rooted. To further conduct a comprehensive fairness evaluation, we additionally select coreset nodes from both extremes of the homophily neighborhood distribution ratio, specifically choosing nodes from the top and bottom $5$ and $10$ of this distribution. For evaluating model performance on the node classification task, we use classification accuracy ACC and the area under the receiver operating characteristic curve AUC.
Figure 4. ACC and AUC comparison across datasets
Figure 5. EO and SP fairness metrics comparison across datasets
Implementation Setup: For a fair comparison with other baseline methods, we employ a standard Graph Convolutional Network (GCN) (Kipf and Welling, 2016) as our encoder. This choice emphasizes our focus on the overall framework. Consistency is maintained across all baselines, which also utilize GCN layers.We implement two-layer GCNs for all datasets and incorporate a two-layer Multilayer Perceptron (MLP) for the predictor network. Our model is implemented using PyTorch and PyTorch Geometric, leveraging their efficient graph processing capabilities. The baseline implementations were adopted from the official PyGDebias (Dong et al., 2023) graph fairness library. Each method was run for three iterations, using the best hyperparameters reported in the original paper. We Perform for $400$ epochs with a learning rate $0.001$ on each dataset on the node classification task for three iterations. All of our experiments were conducted on 16 GB NVIDIA V100 GPUs.
Fairness Paradoxes: To address RQ1, we evaluate the effectiveness of current debiasing techniques for GNNs at the community level and compare the consistency of their debiasing performance when applied to the entire graph. We first measure the group fairness of the GCN model with the debiasing method applied to the original input graph $G$. Next, we measure the group fairness of the same GCN model with the debiasing method applied at the community level. We use the group fairness metrics $\Delta\text{SP}$ and $\Delta\text{EO}$. We observe that the fairness performance across the communities is not as consistent as the fairness score reported for the entire graph across all three real-world datasets for all methods. For example, in Table 3, we observe that the unfairness estimation for the Pokec-z dataset using $\Delta\text{SP}$ and $\Delta\text{EO}$ is higher for communities {2, 3, 4, 5} compared to $\Delta\text{SP}$ and $\Delta\text{EO}$ for the entire dataset when FairGNN debiasing method is used. We also observed Simpson’s paradox on Pokec-z with UGE: the accuracy of Group_00 (66.06) is higher than Group_1111 (62.36) on the entire graph, but on the community level, for 4 out of 5 communities, predictions on Group_1111 have better accuracy.
[…]
4.2. Model Analysis
Figure 6. Comparison of $\Delta$EO and $\Delta$SP across different coreset selection approaches. Comm indicates community.
Ablation Study: To Validate the coreset section of ComFairGNN, we consider different variants of selection approaches based on their homophily neighborhood distribution, section  3.2. (1) Random node selection: we select an equal number of $n$ coreset nodes from groups $S_{0}$ and $S_{1}$ of the entire train set. (2) Top $n$ homophily neighborhood ratio distribution: we select the top $n$ highly homophily nodes of both groups $S_{0}$ and $S_{1}$ from each community. (3) we select $n$ number of coreset nodes of both groups $S_{0}$ and $S_{1}$ from each community randomly. We report the result in figure 6 and make several observations. Firstly, without selecting the $n$ coreset nodes from both extremes of homophily ratio distribution, we often get worse fairness performance, which means representative nodes from both high and low homophily ratio distribution can mitigate the structural bias at the community level. Secondly, by selecting the $n$ coreset nodes randomly from the entire train set, the fairness metric generally becomes worse. Thus, selecting the coreset at the community level is effective in driving the debiasing context at the local subgraph
5. Related Work
5.1. Fairness in Graph
Fairness in Graph Neural Networks (GNNs) is an emerging area of research, researchers have performed signifi- cant work to measure the fairness of those systems(Wang et al., 2019; Fan et al., 2021; li2021user; Kleindessner et al., 2019; Horowitz and Kamvar, 2010). Generally, fairness can be categorized into group fairness and individual fairness. Group fairness focuses on ensuring that the model does not make decisions that favor or disadvantage certain groups. Group fairness is one of the mainly studied. EDITS (Dong et al., 2022) modifies the adjacency matrix and node attributes to produce unbiased graph-structured data by reducing the Wasserstein distance between different groups. Additionally, some studies (Dai and Wang, 2021; Wang et al., 2022b) achieve fairness in GNNs through adversarial training, resulting in a model that generates node embeddings indistinguishable with respect to the sensitive attribute. FairGNN (Dai and Wang, 2021) incorporates fairness regularization to address bias in graph-based tasks. It aims to ensure equitable treatment of different groups in both learned representations and final predictions. UGE (Wang et al., 2022a) addresses the unbiased graph embedding problem by learning node representations from a bias-free underlying graph. This approach eliminates the influence of sensitive node attributes on the embedding process. Another aspect of fairness, often referred to as counterfactual fairness, ensures that changes in sensitive attributes do not impact the prediction results (Agarwal et al., 2021a). NIFTY (Agarwal et al., 2021b) enhances counterfactual fairness and stability in node representations through a novel triplet-based objective function. It employs layer-wise weight normalization using the Lipschitz constant to improve performance and reliability. GEAR addresses graph unfairness through counterfactual graph augmentation and adversarial learning for attribute-insensitive embeddings. These combined approaches effectively reduce bias in graph-based models. GEAR (Ma et al., 2022) addresses graph unfairness through counterfactual graph augmentation and adversarial learning for attribute-insensitive embeddings. These combined approaches effectively reduce bias in graph-based models. Deg-FairGNN (Liu et al., 2023) aims to ensure fair outcomes for nodes with varying degrees within a group. Despite they introduce attribute and structural group fairness, overlooked the inherent local structural bias. In terms of individual fairness, nodes sharing similar attributes are treated equivalently (Dong et al., 2021). For example, REDRESS (Dong et al., 2021) proposes a ranking-based individual fairness optimization method to encourage each individual node relative similarity ranking list to the other nodes during the input and output. InFoRM (Kang et al., 2020) assesses GNN individual fairness using node similarity matrices. It incorporates a modified (d1, d2)-Lipschitz property to address fairness challenges in interconnected graph structures, moving beyond node features alone. However, none of these works are designed for community-level fairness on graphs.
5.2. Detecting Fairness Paradox
There have been prior works that study the fairness paradoxes due to historical bias using statistical measures of fairness and counterfactual fairness. Counterfactual fairness (Kusner et al., 2017) introduces a path-specific causal effect on the prediction outcomes in discrimination analysis. The study demonstrates the paradox that arises when removing the sensitive attribute to achieve fairness; elements of the input feature can still contain discriminatory information that may not be immediately apparent. (Prost et al., 2022) demonstrates Simpson’s Paradox in the context of recommender system fairness, showing that the aggregate metric of fairness is influenced by differences in user distribution across the subgroups.
6. Conclusion
In this paper, we investigate the effectiveness of existing graph debiasing methods and the fairness evaluation paradoxes at the community level. We first group the nodes of the graph into communities based on their structural similarities, then assess the fairness of GNNs with debiasing methods for different demographic subgroups. Our findings demonstrate fairness inconsistencies and paradoxes across various graph communities. By removing the absolute values in $\Delta\text{SP}$ and $\Delta\text{EO}$, we show that absolute measures might dilute these effects, leading to an underestimation of bias in specific communities within the large graph. To address the community-level bias inherent in GNN neighborhood aggregation, we introduce ComFairGNN, a novel fairness-aware GNN framework compatible with various neighborhood aggregation-based GNN architectures. Our approach targets local structural bias by modulating representative coreset nodes from diverse communities through embedding similarity contrast. Comprehensive experiments on three benchmark datasets demonstrate ComFairGNN’s efficacy, yielding promising results in both accuracy and fairness metrics.

### v2

@@ 3.3 Training Objective and Constraints:
$\mathcal{L}_{task}=-\frac{1}{|Y|}\sum_{i}\sum_{j}Y_{i,j}\log(\hat{Y}_{i,j})$ (7)
@@
$\mathcal{L}_{fair}=\left|\frac{1}{|c_{y=0}|}\sum\limits_{i,j\in c_{y=0}}\text{sim}(x_{i},x_{j})-\frac{1}{|c_{y=1}|}\sum\limits_{i,j\in c_{y=1}}\text{sim}(x_{i},x_{j})\right|$ (8)
@@
$\mathcal{L}=\mathcal{L}_{\text{task}}+\mathcal{L}_{\text{fair}}$ (9)
4 Experiments
@@
Table 1: Comparison of GNNs on utility and bias at the graph level with different debiasing methods.
Dataset Nodes Edges Features Classes Comm_1 Comm_2 Comm_3 Comm_4 Comm_5
Pokec-z 67,796 882,765 276 2 309 282 259 1412 302
Pokec-n 66,569 198,353 265 2 428 360 361 389 -
Facebook 1,034 26,749 573 2 55 37 65 - -
Table 2: Dataset statistics and community distribution
4.1 Experimental Setup and Datasets
Datasets. We evaluate on three real-world datasets for node classification: Pokec-z, Pokec-n, and Facebook. The Pokec datasets are subsets from Slovakia’s Pokec social network [19], where nodes represent users, edges represent friendships, and location is the sensitive attribute for classifying users’ work fields. The Facebook dataset [11] similarly uses friendship connections with gender as the sensitive attribute. Dataset statistics are summarized in Table 2.
Evaluation Metrics We use ACC and AUC scores to evaluate the utility performance. To evaluate fairness, we use two commonly used fairness metrics, i.e., $\Delta_{DP}=|P(\hat{y}=1|s=0)-P(\hat{y}=1|s=1)|$ [7] and $\Delta_{EO}=|P(\hat{y}=1|y=1,s=0)-P(\hat{y}=1|y=1,s=1)|$ [10]. $\hat{y}$ and $y$ denote the node label prediction and ground truth, respectively. For $\Delta_{DP}$ and $\Delta_{EO}$, a smaller value indicates a fairer model prediction.
Baseline Methods: We investigate the effectiveness of four state-of-the-art GNN debiasing methods at a graph level and community level, namely FairGNN [4], NIFTY [1], and UGE [22]. FairGNN utilizes adversarial training to eliminate sensitive attribute information from node embeddings. NIFTY improves fairness by aligning predictions based on both perturbed and unperturbed sensitive attributes. This approach generates graph counterfactuals by inverting the sensitive feature values for all nodes while maintaining all other attributes. UGE derives node embeddings from a neutral graph that is free of biases and unaffected by sensitive node attributes, addressing the issue of unbiased graph embedding. We use GCN [12] as the backbone GNN model.
Implementation Setup: We use a two-layer GCN [12] encoder and MLP predictor for all experiments, consistent with baseline methods implemented from PyGDebias [6]. Each method runs for 400 epochs with learning rates in {0.1,0.01,0.001}, using hyperparameters from their original papers and repeating three times. Experiments were conducted on 16GB NVIDIA V100 GPUs.
(a) Performance at a graph level
(b) Performance on Pokec-z communities
(c) Performance on Pokec-n communities
(d) Performance on Fb communities
Figure 4: Performance comparison of different GNN debiasing methods across datasets and communities. (a) Graph-level performance across different datasets (Pokec-z, Pokec-n, Facebook). (b, c, and d) Community-wise performance breakdown on Pokec-z, Pokec-n and Facebook datasets. Showing AUC(%) and EO(%) metrics.
Fairness Paradoxes: To address RQ1, we evaluate the effectiveness of current debiasing techniques for GNNs at the community level and compare the consistency of their debiasing performance when applied to the entire graph. We first measure the group fairness of the GCN model with the debiasing method applied to the entire original input graph $G$. Next, we measure the group fairness of the same GCN model with the debiasing method applied at the community level. We observe that the fairness performance across the communities is not as consistent as the fairness score reported for the entire graph across datasets for all baseline methods. For example, in figure 4 {b, c, d}, we observe that the unfairness estimation for Pokec-z dataset using $\Delta\text{EO}$ is higher for communities {2, 3, 4, 5} compared to $\Delta\text{EO}$ for the entire dataset when FairGNN debiasing method is used.
Over-Simplification of Fairness Evaluation RQ2: The absolute values in $\Delta\text{SP}$ and $\Delta\text{EO}$ metrics can mask local disparities across communities. While overall metrics might show moderate fairness when averaged, they obscure cases where a group experiences high bias in some communities and low bias in others, as shown in Figure 4. This averaging effect fails to capture the actual fairness experienced by individuals in different parts of the graph.
Evaluation and Main Results (RQ3) As GNNs learn by neighborhood aggregation section 2.1, we evaluate fairness within different subgraphs or communities within the larger graph. This can help identify and address local disparities more effectively. We select $S_{0}$ and $S_{1}$ nodes within the community to evaluate the model’s disparity locally, which may present more prominent biases. We select top and bottom $15$ sample homophily neighborhood distribution ratio section 3.2, from each community to formulate the representative coreset nodes from the training set during training. These nodes are sufficient to cover the local contextual structures where biases are rooted. As shown in Table  1, ComFairGNN has both the best overall classification performance than the baselines methods. In terms of fairness, ComFairGNN respectively reduces $\Delta\text{SP}$ and $\Delta\text{EO}$ by 4.9% and 4.1%, compared with the best performed baseline on both Pokec-z and Pokec-n datasets. Additionally, as shown in Figure 4, this pattern is consistent across all communities in the datasets. Hence, the fairness improvement achieved by ComFairGNN aligning with our motivation.
Ablation Study: To validate ComFairGNN’s, we investigate different coreset sizes of $C$ hyperparameter. As shown in figure 5 $C=30$ sample size relatively performs better.
Figure 5: AUC and EO Performance Comparison Across Different Coreset Sizes $C$
5 Related Work
5.1 Fairness in Graph
Research in Graph Neural Networks (GNNs) fairness has produced several approaches to measure and mitigate bias. Recent works have primarily focused on group fairness, ensuring models don’t discriminate against certain groups. EDITS [5] modifies the adjacency matrix and node attributes to reduce the Wasserstein distance between different groups. FairGNN [4] incorporates fairness regularization to ensure equitable treatment in both representations and predictions. UGE [22] learns unbiased node representations by eliminating sensitive attribute influences, while NIFTY [2] enhances counterfactual fairness through triplet-based objectives and layer-wise weight normalization. DegFairGNN [13] addresses fairness for nodes with varying degrees within groups. However, this approach is computationally expensive for large graphs. Additionally, like other existing methods, DegFairGNN has not fully addressed the inherent local structural bias in graphs.
6 Conclusion
This paper examines fairness paradoxes in graph debiasing methods at the community level. By analyzing structural communities, we reveal fairness inconsistencies across demographic subgroups that are masked when using absolute measures of $\Delta\text{SP}$ and $\Delta\text{EO}$. To address community-level bias in GNN neighborhood aggregation, we propose ComFairGNN, a fairness-aware framework that modulates representative coreset nodes through embedding similarity contrast. Experiments on three benchmark datasets demonstrate that ComFairGNN effectively improves both accuracy and fairness metrics.
Acknowledgments
This work is partially supported by the National Science Foundation under Grant No. 2152117 and NSF-CAREER 2237831. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.